# Non-Asymptotic Performance Analysis of DOA Estimation Based on Real-Valued Root-MUSIC

### v2

@@ 3.1 Deviation of the true DOAs
$\begin{split}\left.\frac{\partial f(z,\tilde{r})}{\partial z}\right|_{z=r_{k}}&=A[K_{t}(r_{k})^{{}^{\prime}}[K_{m}(r_{k})G(r_{k})]\\ &\quad\left.+K(r_{k})[K_{m}(r_{k})G(r_{k})]^{{}^{\prime}}]\right|_{\tilde{r}=r_{k}+\Delta r_{k}}\\ &=A[K_{t}(r_{k})^{{}^{\prime}}K_{m}(r_{k})G(r_{k})]\\ &\doteq 2jAr_{k}^{*}Im(\frac{\Delta r_{k}}{r_{k}})K_{m}(r_{k})G(r_{k})\end{split}$ (16)
@@ 3.2 Deviation of the mirror DOAs
$\begin{split}\left.\frac{\partial f(z,\tilde{r})}{\partial z}\right|_{z=r_{k}^{*}}&=A[K_{m}(r_{k}^{*})^{{}^{\prime}}[K_{t}(r_{k}^{*})G(r_{k}^{*})]\\ &\quad\left.+K_{m}(r_{k}^{*})[K_{t}(r_{k}^{*})G(r_{k}^{*})]^{{}^{\prime}}]\right|_{\tilde{r}=r_{k}^{*}+\Delta(r_{k}^{*})}\\ &=\left.A[K_{m}(r_{k}^{*})^{{}^{\prime}}K_{t}(r_{k}^{*})G(r_{k}^{*})]\right|_{\tilde{r}=r_{k}^{*}+\Delta(r_{k}^{*})}\\ &\doteq 2jAr_{k}Im(\frac{\Delta r_{k}^{*}}{r_{k}^{*}})K_{t}(r_{k}^{*})G(r_{k}^{*})\end{split}$ (22)

### v3

@@ 3.1 Deviation of the true DOAs
$\begin{split}\left.\frac{\partial f(z,\tilde{r})}{\partial z}\right|_{z=r_{k}}&=A[K_{t}(r_{k})^{{}^{\prime}}[K_{m}(r_{k})G(r_{k})]\\ &\qquad\left.+K(r_{k})[K_{m}(r_{k})G(r_{k})]^{{}^{\prime}}]\right|_{\tilde{r}=r_{k}+\Delta r_{k}}\\ &=A[K_{t}(r_{k})^{{}^{\prime}}K_{m}(r_{k})G(r_{k})]\\ &\doteq 2jAr_{k}^{*}Im(\frac{\Delta r_{k}}{r_{k}})K_{m}(r_{k})G(r_{k})\end{split}$ (16)
@@ 3.2 Deviation of the mirror DOAs
$\begin{split}\left.\frac{\partial f(z,\tilde{r})}{\partial z}\right|_{z=r_{k}^{*}}&=A[K_{m}(r_{k}^{*})^{{}^{\prime}}[K_{t}(r_{k}^{*})G(r_{k}^{*})]\\ &\qquad\left.+K_{m}(r_{k}^{*})[K_{t}(r_{k}^{*})G(r_{k}^{*})]^{{}^{\prime}}]\right|_{\tilde{r}=r_{k}^{*}+\Delta(r_{k}^{*})}\\ &=\left.A[K_{m}(r_{k}^{*})^{{}^{\prime}}K_{t}(r_{k}^{*})G(r_{k}^{*})]\right|_{\tilde{r}=r_{k}^{*}+\Delta(r_{k}^{*})}\\ &\doteq 2jAr_{k}Im(\frac{\Delta r_{k}^{*}}{r_{k}^{*}})K_{t}(r_{k}^{*})G(r_{k}^{*})\end{split}$ (22)